OpenStudy (anonymous):

What is the exact value for the expression the square root of 153. − the square root of 17. + the square root of 68.? Simplify if possible. the square root of 17. 4 the square root of 17. 2 the square root of 51. 8 the square root of 51.

OpenStudy (anonymous):

\[3\sqrt{17} - \sqrt{17} + 2\sqrt{17}\] the answer is\[4\sqrt{17}\]

OpenStudy (anonymous):

Are you 100%?